OpenStudy (anonymous):

Find an expression for a rational function f(x) that satisfies the conditions: a horizontal asymptote at y = 0, vertical asymptote at x = 0, and contains the point (1,11)

OpenStudy (anonymous):

All of the above f(x) = 11/x f(x) = 11/x^3 f(x) 11/x^2

OpenStudy (john_es):

All of the above satisfy the conditions.